What is the frequency of a photon with a wavelength of 159 nm
1
Expert's answer
2020-02-25T05:02:56-0500

The frequency of a photon is calculated by the formula:

"\lambda=\tfrac{c}{\nu}" , where "\u0441=3*10^8" m/s.

A wavelength in nm translate to m:

159 nm = "1.59*10^{-7}" m.

"\lambda=\tfrac{3*10^8}{1.59*10^{-7}}=1.88*10^{15}" Hz.


Answer: "1.88*10^{15}" Hz.